Divers fall from a point 36 meters above the sea.
a) How long does it take them to hit the water?


b) What is their velocity when they enter the water?



I dont know how to solve

Answer by scott8148(6628)

gravitational acceleration is 9.8m/s^2 (meters per second per second)

starting from rest , distance equals one half acceleration times time squared ___ d = .5 a t^2

36 = .5 * 9.8 * t^2 ___ 72/9.8 = t^2 ___ use your calculator to find t

velocity equals acceleration times time ___ v = at = 9.8 * t ___ use the calculated value of t to find the velocity